OpenStudy (anonymous):

absolute value x + 1

OpenStudy (anonymous):

\[\left| x +1 \right|\]

OpenStudy (anonymous):

why are u cursing at me?

OpenStudy (anonymous):

wat do u want to do with this?

OpenStudy (anonymous):

i have to write with absolute value so is the answer just x+1

OpenStudy (amistre64):

carnak say ...... maybe

OpenStudy (anonymous):

huh?

OpenStudy (amistre64):

yep .... thats how we feel when you expect us to read your mind. Carnak is an old skit by johnny carson where he plays a psychic.

OpenStudy (anonymous):

sorry i meant i have to write without absolute value

OpenStudy (amistre64):

:) if all you have to do is remove the |...| then yeah, thats it. but its still hard to figure out what it is that you need for a good answer

OpenStudy (amistre64):

absolute values tend to be piecewise defied functions; so i dont know if simply removing the bars is a good enough response

OpenStudy (anonymous):

ok then what about this question: Solve the inequality \[\left| x + 5 \right|\ge2\]

OpenStudy (amistre64):

\[|x+1|=\left[\begin{array}c x+1&x\ge-1\\-x-1&x\lt -1\end{array}\right]\] but i cant recall if that second part is good

OpenStudy (anonymous):

WHOA OK!

OpenStudy (amistre64):

|x+5| >/ 2 x+5 >/ 2 and -x-5 >/ 2 x >/ -3 -x >/ 7 x>/ -3 x </ -7

OpenStudy (anonymous):

so i make the x+5 negative and not the 2 on the outside negative

OpenStudy (anonymous):

solution for |x+5|≥2 is below. if |x|>a then x<−a or x>a so for \[\left| x+5 \right| \ge2\]there are two cases \[x+5\le−2\]which implies\[x\le−7\]and \[x+5\ge2\]which implies\[x\ge−3\]

OpenStudy (anonymous):

then the solution set \[(-\infty,-7]\cup[-3,\infty)\]

OpenStudy (anonymous):

for \[\left| x+1 \right |\] \[\left| x+1 \right|= \left\{\begin{matrix} -x-1 & x<-1\\ x+1 & x>-1 \end{matrix}\right. \]